OpenStudy (anonymous):

In 1977, the price of a scientific calculator was $175. In 2004, the price was $15. Assuming the change in price was linear, what was the approximate price of a scientific calculator in 1998? $23.00 $27.00 $51.00 $60.00

OpenStudy (anonymous):

Is 1977 year 0 or year 1?

OpenStudy (anonymous):

51

OpenStudy (anonymous):

1977 is year 0. since it's Initial value.

OpenStudy (anonymous):

omg i love you guyss, (: thanks so much, 25 more questions all most done (:

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Bailey Nae try solving them your self at first. otherwise you will not gett anything.

OpenStudy (anonymous):

i do, i try them 1st Lol i just keep gettin the answers wrong and call me nae hun (:

OpenStudy (anonymous):

okay this is a y of 2 minus y of 1 over x of 2 - x of 1 type of problem. Solve the equation of the line. Your x value is the year and the y value is the $$. The two points on the line are (0,175) and then (27, 15) Does this make sense so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

175 - 15 = 160 0-27 = 27 160/27 is the slope of the line (M in the equation y =mx+b)

OpenStudy (anonymous):

Since you have the y values, and the slope all you need to do now is solve for the x values

OpenStudy (anonymous):

so its 27?

OpenStudy (anonymous):

160/27 is the slope of the line in the equation y=mx+b. you started with two points on the line and needed to find the equation of the line so that you can solve for the rest of the questions. So step one is to find the equation of the line y=160/27x + b Now, plug in the y values one at a time and solve for X.

OpenStudy (anonymous):

Taking a look at the first one ($23.00) y= $23.00 23 = 160/27x + b

OpenStudy (anonymous):

so it 23?????????? im confused

OpenStudy (anonymous):

Okay... how familiar are you with solving an equation of a line given two points on the line? I will explain step by step but need to know where to begin

OpenStudy (anonymous):

http://www.purplemath.com/modules/slopgrph.htm